OpenStudy (richyw):

Can you just straight up type latex in? I don't like the equation editor. Especially because I can't seem to find how to put in fractions.

OpenStudy (lgbasallote):

yep you can

OpenStudy (lgbasallote):

for example: \[\text{\[\LaTeX\]}\]

OpenStudy (lgbasallote):

if you type that you will get \(\LaTeX\)

hero (hero):

\(\frac{a}{b}

hero (hero):

enclose with \)

hero (hero):

\(\frac{a}{b}\)

OpenStudy (lgbasallote):

\[\text{\[{a \over b}\]} \rightarrow {a \over b}\]

hero (hero):

I can't stand equation editor

OpenStudy (richyw):

\[\frac{1}{4\pi \epsilon_\circ}\]

OpenStudy (richyw):

awesome, that is going to save me SO MUCH time! thanks.

OpenStudy (lgbasallote):

\[\text {\[\frac ab\]} \implies \frac ab\]

hero (hero):

lgba is showing off :P

OpenStudy (lgbasallote):

lol just showing the variations *whistle*

hero (hero):

You can use the format I showed you to write fractions like \(\large\frac{5\pi}{2p_o}\) in text.

OpenStudy (lgbasallote):

you can also do textception \[\large \text{\[\text{\text{\text{\text}}}\]}\]

OpenStudy (richyw):

so I put \(\frac{a}{b}\) for inline and if I want to display I put \[\frac{a}{b}\]. The live preview is also awesome

hero (hero):

I think you have it all figured out now.

hero (hero):

If you want to know more advanced features, talk to lgba or Parth.